Recall the example of rolling a six-sided die. This is an example of a discrete uniform random variable, so named because the probability of observing each distinct outcome is the same, or uniform, for all outcomes. Let Y be the discrete uniform random variable that equals the face-value after a roll of an eight-sided die. (The die has eight faces, each with number 1 through 8.) Calculate E(Y ), Var(Y ), and StdDev(Y )


1
Expert's answer
2021-02-24T07:02:08-0500

We have that sample space is "S=\{1,2,3,4,5,6,7,8\}"

The die is fair thus each of the eight faces has an equally likely probability of occurring, i.e., 1/8.

The expected value is calculated by the formula:

"E(Y) =\sum yP(y)"

"(1+2+3+4+5+6+7+8)\cdot\frac{1}{8}=4.5"

Variation is calculated by the formula:

"var(Y)=E(Y-E(Y)^2)=\sum(y-E(Y))^2P(y)=\sum(y-E(Y))^2\cdot\frac{1}{8}"

"((1-4.5)^2+(2-4.5)^2+(3-4.5)^2+(4-4.5)^2+(5-4.5)^2+(6-4.5)^2+(7-4.5)^2+(8-4.5)^2)\cdot\frac{1}{8}=5.25"

Standard deviation is calculated by the formula:

"stddev(Y)=\sqrt{var(Y)}=\sqrt{5.25}=2.29"


Answer:

E(Y) = 4.5

var(Y) = 5.25

steddev(Y) = 2.29